OpenStudy (anonymous):

4y^4+y^2+2-8y^2+2y^4 this is simplifying polynominals i need help

OpenStudy (anonymous):

4y^4+y^2+2-8y^2+2y^4 group y^4 terms and group y^2 terms =(4y^4+2y^4)+(y^2-8y^2)+2 =(6 y^4 )+(-6 y^2) +2 =6 y^4 -6 y^2 +2